OpenStudy (anonymous):

1. Find the slope of line passing through the points (4,8) and (6, -2). A. 1/5 B. 5 C. -1/5 D. -5

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Use the point slope formula: \[m (slope ) = (y_2 - y_1) / (x_2 - x_1)\]

OpenStudy (anonymous):

So (-2 - 8) / (6 - 4) = -10 / 2 = -5

OpenStudy (anonymous):

thnk u